Vapour pressure of chloroform (CHCl3) and dichloromethane (CH2Cl2) at 25C are 200mm Hg and 41.5 mm Hg respectively. Vapour pressure of the solution obtained by mixing 25.5g of (CHCl2) and 40g of (CH2Cl2) at the same temperature will be 
(molecular mass of CHCl3 = 119.5 u and molecular mass of CH2Cl2 = 85 u ) 

Grade:12

1 Answers

Arun
6 years ago

Molar mass of CH2Cl2 = 12 × 1 + 1 × 2 + 35.5 × 2 = 85 g mol–1

   Molar mass of CHCl3 = 12 × 1 + 1 × 1 + 35.5 × 3 = 119.5 g mol-1

Moles of CH2Cl2= 40 g /85 g mol = 0.47 mol

Moles of CHCl3 = 25.5 g /119.5 g mol = 0.213 mol

Total number of moles = 0.47 + 0.213 = 0.683 mol

 

Molefraction of component 2

 = 0.47 mol / 0.683 mol = 0.688

Molefraction of component 1

= 1.00 – 0.688 = 0.312

 

We know that

Ptotal = (p2-p1)*x

  = 200 + (415 – 200) × 0.688

  = 200 + 147.9

  = 347.9 mm Hg